OpenStudy (anonymous):

find the slope between (8,0) and (2,5)

OpenStudy (anonymous):

Slope formula m=\[\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (anonymous):

slop=y2-y1/x2-x1 slop =5-0/2-8 ==-5/6

OpenStudy (anonymous):